OpenStudy (carolineclayton):

I'm having trouble figuring this out.. sqrt(3x^7y^-9)*sqrt(6x^8y^13) Thank you so much!

OpenStudy (xapproachesinfinity):

what do you want to figure out !

OpenStudy (xapproachesinfinity):

use property \(\sqrt{a}\times \sqrt{b}=\sqrt{a\times b}\)

OpenStudy (xapproachesinfinity):

so in this care a is 3x^7y^-9 and b is 6x^8y^13

OpenStudy (carolineclayton):

I need to turn it into the form \[a \sqrt{b}\]

OpenStudy (xapproachesinfinity):

\[\sqrt{3x^7y^{-9}}\sqrt{6x^8y^{13}}=\sqrt{3x^7y^{-9}6x^8y^{13}}\]

OpenStudy (xapproachesinfinity):

simplify

OpenStudy (carolineclayton):

\[\sqrt{9x^(15)y^4}\]

OpenStudy (carolineclayton):

In other words, sqrt(9x^15v^4)

OpenStudy (xapproachesinfinity):

yes that is good now reduce it

OpenStudy (xapproachesinfinity):

y^4 and 9 can be reduced to 3y^2 so you get 3y^2sqrt{x^15}

OpenStudy (xapproachesinfinity):

actually where did you get 9 it should be 18 not 9

OpenStudy (xapproachesinfinity):

but you rewrite it like 2x9 so the actual result is 3y^2sqrt{2x^15}

OpenStudy (carolineclayton):

So the final answer was 3x^7y^2sqrt(2x) And it's correct! Thanks so much!

OpenStudy (xapproachesinfinity):

yeah you could write that way

OpenStudy (xapproachesinfinity):

what you need to know is that \[a^{p/q}=\sqrt[q]{a^p}\]

OpenStudy (xapproachesinfinity):

it will make your life easier

OpenStudy (xapproachesinfinity):

that is by definition of rational powers

OpenStudy (carolineclayton):

Ok, great! Thank you!